An economy’s marginal propensity to consume out of disposable income is 0.75, its
marginal and average tax rate is 0.15, and its marginal and average propensity to import
is 0.15. What is the increase in government expenditure needed to prevent a fall in
equilibrium income of 20 billion Dollar?

Hi experts.,,
Please help me as fast as possiple because i want the answer urgent.
1
Expert's answer
2014-05-21T08:31:43-0400
c = 0.75, t = 0.15, im = 0.15. Find increase in government expenditure (E) needed to prevent a fall in equilibrium GDP of $20
billion.

Change in E = change in GDP/Me
Expenditure multiplier Me = 1/(1-(c-im)(1-t)) = 1/(1-(0.75-0.15)(1-0.15)) =1/0.49 = 2.04
Change in E = 20/2.04 = $9.8 billion